# Can one size fit all?: Measuring Failure in Multi-Document Summarization Domain Transfer

### v1

@@ Appendix C Raw Results
Table 4: Statistics for each dataset used in this work: Multi-News, Multi-XScience, and ConvoSumm-Reddit. The number of tokens in the (ground truth) summary and documents are measured with the BART tokenizer. Since we remove examples with missing information, our numbers might not match those in the original dataset papers (Appendix A).
Domain Factuality Accuracy Quality
Train/Test Cons. (UE) Rel. (UE) RL-Sum MoverScore Fluency (UE) Coh. (UE)
ConvoSumm Convo 0.64 0.70 0.21 0.06 0.91 0.60
News 0.64 0.62 0.19 0.17 0.89 0.66
Science 0.67 0.55 0.15 0.01 0.69 0.46
All 0.65 0.62 0.18 0.08 0.83 0.57
PRIMERA Convo 0.65 0.73 0.21 0.09 0.92 0.71
News 0.74 0.82 0.25 0.21 0.93 0.78
Science 0.96 0.68 0.14 0.05 0.83 0.86
All 0.78 0.74 0.20 0.12 0.89 0.78
SummN Convo 0.65 0.63 0.21 0.07 0.89 0.62
News 0.69 0.78 0.21 0.20 0.89 0.80
Science 0.65 0.39 0.15 -0.02 0.66 0.27
All 0.66 0.60 0.19 0.08 0.81 0.56
Llama 3.1 Convo 0.80 0.92 0.14 0.07 0.93 0.94
News 0.82 0.90 0.19 0.19 0.94 0.93
Science 0.85 0.90 0.12 0.08 0.94 0.95
All 0.83 0.91 0.13 0.08 0.93 0.95
@@
Table 5: Multi-document summarization (MDS) model performance when trained and tested on the same domain. The results for “All” are averaged across all domains for each model. Results are on the test splits of each dataset. Metrics followed by (UE) are from UniEval. For all metrics a greater score is better.
Domain Factuality Accuracy Quality
Train Test Fact. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
ConvoSumm Convo News 0.05 -0.41 -0.37 -0.44 -1.67 0.00 -0.25
Convo Science 0.06 -0.09 -0.25 -0.33 -0.72 -0.01 0.02
News Convo -0.11 0.30 -0.46 -0.31 -1.12 0.02 0.19
News Science -0.07 0.10 -0.41 -0.28 -0.90 0.05 0.02
Science Convo -0.07 -0.30 -0.16 -0.10 -2.79 0.03 -0.13
Science News -0.01 -0.28 -0.01 -0.08 -1.90 0.07 -0.07
All 0.01 -0.02 -0.11 -0.28 -0.26 -1.52 0.03 -0.04
PRIMERA Convo News 0.20 -0.31 -0.05 -0.22 -0.22 -0.09 -0.08
Convo Science 0.48 -0.07 -0.24 -0.36 -0.47 -0.10 0.20
News Convo -0.22 0.00 -0.61 -0.57 -1.03 -0.07 0.00
News Science 0.29 -0.17 -0.48 -0.45 -0.78 -0.11 0.09
Science Convo -0.47 -0.12 -0.20 -0.13 -0.74 -0.27 -0.25
Science News -0.11 0.04 0.43 0.30 2.25 -0.21 -0.03
All 0.03 -0.10 -0.19 -0.24 -0.16 -0.14 -0.01
SummN Convo News 0.31 -0.08 -0.22 -0.30 -0.67 0.04 0.14
Convo Science 0.26 -0.02 -0.31 -0.35 -0.79 0.04 0.31
News Convo -0.32 0.05 -0.69 -0.60 -1.22 -0.00 -0.00
News Science -0.17 -0.21 -0.52 -0.41 -1.00 -0.01 -0.25
Science Convo -0.00 -0.34 -0.59 -0.51 5.56 0.05 0.14
Science News 0.17 0.03 0.37 0.17 -3.74 0.26 0.72
All -0.06 0.04 -0.09 -0.32 -0.33 -0.31 0.07 0.18
Llama 3.1 Convo News 0.04 -0.04 0.69 0.42 1.74 0.01 -0.01
Convo Science 0.07 -0.03 0.05 0.03 0.29 -0.00 0.03
News Convo -0.02 0.04 -0.48 -0.36 -0.72 -0.01 0.03
News Science 0.05 -0.00 -0.42 -0.30 -0.56 -0.01 0.04
Science Convo -0.06 0.02 -0.10 -0.05 -0.34 -0.01 -0.02
Science News -0.03 -0.05 0.70 0.49 1.11 0.01 -0.05
All 0.01 -0.01 0.07 0.04 0.25 -0.00 0.00
Llama 3.1
(0-shot)
zero-shot Convo -0.01 -0.03 -0.16 -0.07 -0.23 -0.01 -0.02
zero-shot News 0.02 -0.07 -0.01 0.06 -0.01 -0.00 -0.03
zero-shot Science 0.01 -0.01 0.01 0.02 0.01 -0.01 0.01
All 0.01 -0.04 -0.05 0.00 -0.07 -0.01 -0.01
Table 6: Multi-document summarization (MDS) model relative performance when trained and tested on different domains. Results are on the test splits of each dataset. A positive number shows a performance gain and a negative number is a performance loss as compared to the model’s scores on the in-domain data. A relative change of over $1$ is possible with MoverScore because of its $[-1,1]$ range.
Domain Factuality Accuracy Quality
Train Test Fact. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
Convo News 0.15 -0.21 0.01 -0.14 -0.20 -0.01 -0.05
Science 0.22 -0.05 -0.19 -0.25 -0.42 -0.02 0.14
All 0.18 -0.13 -0.09 -0.19 -0.31 -0.01 0.04
News Convo -0.17 0.10 -0.56 -0.46 -1.02 -0.02 0.05
Science 0.03 -0.07 -0.46 -0.36 -0.81 -0.02 -0.03
All -0.07 0.02 -0.51 -0.41 -0.92 -0.02 0.01
Science Convo -0.15 -0.18 -0.26 -0.20 0.42 -0.05 -0.06
News 0.01 -0.06 0.37 0.22 -0.57 0.03 0.14
All -0.07 -0.12 0.06 0.01 -0.07 -0.01 0.04
zero-shot Convo -0.01 -0.03 -0.16 -0.07 -0.23 -0.01 -0.02
News 0.02 -0.07 -0.01 0.06 -0.01 -0.00 -0.03
Science 0.01 -0.01 0.01 0.02 0.01 -0.01 0.01
All 0.01 -0.04 -0.05 0.00 -0.07 -0.01 -0.01
Table 7: Multi-document summarization (MDS) model relative performance when trained and tested on different domains. The results are averaged across all model types. Results are on the test splits of each dataset.
Domain Factuality Accuracy Quality
Train Test Cons. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
ConvoSumm Convo News 0.68 0.41 0.20 0.12 -0.04 0.92 0.45
Convo Science 0.68 0.63 0.23 0.14 0.02 0.91 0.61
News Convo 0.57 0.81 0.21 0.13 -0.02 0.90 0.78
News Science 0.60 0.69 0.24 0.13 0.02 0.93 0.67
Science Convo 0.63 0.38 0.21 0.14 -0.02 0.71 0.40
Science News 0.67 0.39 0.25 0.14 -0.01 0.73 0.43
PRIMERA Convo News 0.78 0.50 0.31 0.16 0.07 0.84 0.65
Convo Science 0.96 0.68 0.25 0.14 0.05 0.83 0.86
News Convo 0.58 0.82 0.18 0.11 -0.01 0.86 0.79
News Science 0.96 0.68 0.25 0.14 0.05 0.83 0.86
Science Convo 0.51 0.60 0.20 0.12 0.01 0.60 0.64
Science News 0.86 0.71 0.35 0.18 0.15 0.65 0.83
SummN Convo News 0.85 0.58 0.24 0.15 0.02 0.93 0.71
Convo Science 0.82 0.62 0.22 0.14 0.01 0.92 0.81
News Convo 0.47 0.82 0.13 0.08 -0.05 0.89 0.80
News Science 0.57 0.62 0.21 0.12 -0.00 0.89 0.60
Science Convo 0.64 0.26 0.10 0.07 -0.16 0.69 0.31
Science News 0.76 0.40 0.33 0.18 0.07 0.83 0.47
Llama 3.1 Convo News 0.83 0.89 0.41 0.22 0.18 0.94 0.93
Convo Science 0.85 0.90 0.26 0.16 0.09 0.93 0.97
News Convo 0.80 0.94 0.21 0.14 0.05 0.93 0.95
News Science 0.86 0.90 0.24 0.15 0.08 0.94 0.96
Science Convo 0.80 0.93 0.21 0.14 0.05 0.93 0.93
Science News 0.82 0.86 0.39 0.23 0.18 0.94 0.91
Llama 3.1 (0-shot) zero-shot Convo 0.79 0.90 0.21 0.15 0.05 0.93 0.92
zero-shot News 0.83 0.84 0.40 0.23 0.19 0.94 0.90
zero-shot Science 0.86 0.90 0.23 0.16 0.08 0.92 0.96
Table 8: Multi-document summarization (MDS) model performance when trained and tested on the different domains. Results are on the test splits of each dataset. Metrics followed by (UE) are from UniEval. For all metrics a greater score is better.
Stage Setting Multi-News ConvoSumm-Reddit Multi-XScience
General
total number of updates 30000 5000 -
warmup updates 500 200 -
learning rate 3e-5 - -
max input sequence 1024 - -
max tokens (in batch) 2048 - -
update frequency 4 - -
patience 2 - -
label smoothing 0.1 - -
dropout 0.1 - -
attention dropout 0.1 - -
weight decay 0.01 - -
optimizer Adam betas=(0.9, 0.999), eps=1e-8 - -
LR scheduler polynomial decay - -
Gradient clipping 0.1 - -
Stage 1
max input sequence 1024 - -
beam 1 - -
length penalty 2 1 -
max length 450 - -
min length 150 0 -
no-repeat-ngram-size 3 0 -
Stage 2
max input sequence 1024 - -
beam 1 - -
length penalty 2 1 -
max length 550 - -
min length 150 0 -
no-repeat-ngram-size 3 0 -
Stage 3
max input sequence 1024 - -
beam 1 - -
length penalty 2 1 -
max length 900 1024 -
min length 400 0 -
no-repeat-ngram-size 3 0 -
Setting Multi-News* ConvoSumm-Reddit Multi-XScience*
max input sequence 4096 4096 4096
total number of steps 25K 5K 20K
warmup steps 2.5K 200 2K
optimizer Adam Adam Adam
learning rate 3e-5 3e-5 3e-5
batch size 16 16 16
seed 0 42 0
label smoothing 0.1 0.1 0.1
max length 1024 1024 1024
min length 0 0 0
beam 1 1 1
Table 10: Training details for PRIMERA on all datasets. Models marked with (*) were trained by (Xiao et al., 2022b) and only available training information is filled in. Weights initialized from BART-Large.
Setting Multi-News ConvoSumm-Reddit* Multi-XScience
max input sequence 2048 2048 2048
max tokens (in batch) 2048 - -
update frequency 1 1 1
total number of updates 30000 200 200
warmup updates 500 20 20
learning rate 3e-5 3e-5 3e-5
label smoothing 0.1 0.1 0.1
dropout 0.1 - -
attention dropout 0.1 - -
weight decay 0.01 - -
optimizer Adam betas=(0.9, 0.999), eps=1e-8 - -
LR scheduler polynomial decay - -
Gradient clipping 0.1 - -
max length 1024 1024 1024
min length 0 0 0
beam 1 1 1
length penalty 1 1 1

### v2

@@ Appendix E Raw Results
Table 2: Statistics for each dataset used in this work: ConvoSumm-Reddit, Multi-News+, and Multi-XScience. The number of tokens in the (ground truth) summary and documents are measured with the BART tokenizer. Since we remove examples with missing information, our numbers might not match those in the original dataset papers (Appendix˜A).
(a) ConvoSumm-Reddit
(b) Multi-News+
(c) Multi-XScience
Figure 7: Pearson correlation between metrics on the test ground-truth (reference) for the evaluated datasets.
(a) ConvoSumm
(b) SummN
(c) Llama 3.1 8B
(d) PRIMERA
Figure 8: Pearson correlation between metrics across all datasets for each model in the domain-transfer setting.
ConvoSumm-Reddit Multi-News+ Multi-XScience
From $\Delta$(%) To $\Delta$(%) Ref. From $\Delta$(%) To $\Delta$(%) Ref. From $\Delta$(%) To $\Delta$(%) Ref.
Factuality Cons. (UE) 0.74 +11% -9% 0.71 0.80 -5% +8% 0.70 0.72 -5% +1% 0.55
Fact. (P) 2.25 +11% -21% 3.07 2.46 -20% +10% 2.84 2.11 -5% -1% 1.82
Cons. (P) 2.42 +5% -18% 3.06 2.62 -21% +11% 3.16 2.09 +1% -7% 1.86
Ground-Truth Similarity Rel. (UE) 0.69 -1% -15% 0.99 0.73 0% -17% 0.98 0.63 -29% +5% 0.99
R1 0.26 +13% -30% -1.00 0.42 -46% +28% -1.00 0.26 +6% -21% -1.00
RLSum 0.17 -1% -22% -1.00 0.21 -37% +11% -1.00 0.15 +3% -20% -1.00
MoverScore 0.05 -238% -98% 1.00 0.17 -80% -110% 1.00 0.05 -36% -107% 1.00
Quality Rel. (P) 2.62 +5% -18% 3.23 2.81 -15% +8% 3.39 2.37 -4% -2% 2.78
Fluency (UE) 0.85 +2% -5% 0.90 0.76 +4% -4% 0.76 0.77 -9% +7% 0.66
Fluency (P) 2.69 +7% -16% 3.30 2.79 -21% +11% 2.75 2.17 +2% -5% 1.53
Coh. (UE) 0.73 +6% -7% 0.73 0.81 -3% -3% 0.75 0.62 -13% +2% 0.43
Coh. (P) 2.60 +6% -19% 3.39 2.89 -24% +12% 3.70 2.17 +2% -7% 1.51
Comp.-Sent 9.01 +45% -72% 12.03 2.74 -44% +19% 2.08 4.07 -42% +18% 6.10
Comp.-Word 7.18 +13% -24% 8.63 2.47 -18% +8% 1.98 4.40 -8% +4% 5.92
Table 3: Multi-document summarization (MDS) model performance and relative performance aggregated by domain when trained and tested on the same domain, and the relative change ($\Delta$) when transferred to other domains (From), and vice-versa (To). A relative change of over 111 is possible with MoverScore because of its [1,1][-1,1][ - 1 , 1 ] range. For all metrics other than Compression, a higher score is better. Metrics measured by UniEval and Prometheus are marked by (UE) and (P), respectively.
Domain Factuality Ground-Truth Similarity Quality
Train Test C(UE) F(P) C(P) R(UE) R1 RLS MS R(P) Fl(UE) Fl(P) Co(UE) Co(P) Cov(P) C.Sent C.Word
ConvoSumm Convo News+ 0.75 1.83 1.90 0.51 0.24 0.14 0.00 2.08 0.91 2.02 0.57 1.94 1.68 -4.31 -7.54
Convo Science 0.67 1.95 1.97 0.61 0.23 0.14 0.02 2.27 0.91 2.22 0.58 2.03 1.76 -5.17 -9.23
News+ Convo 0.68 1.59 1.78 0.75 0.23 0.13 0.01 1.97 0.70 1.72 0.78 1.74 1.77 -3.62 -2.51
News+ Science 0.69 1.77 1.76 0.70 0.23 0.13 0.04 2.13 0.81 1.79 0.71 1.87 1.89 -2.31 -2.32
Science Convo 0.44 1.22 1.26 0.26 0.22 0.15 -0.01 1.34 0.60 1.27 0.26 1.29 1.25 -9.32 -7.78
Science News+ 0.66 1.81 1.87 0.26 0.30 0.16 0.04 2.10 0.64 1.65 0.30 1.70 1.69 -3.37 -3.99
PRIMERA Convo News+ 0.83 1.97 2.08 0.41 0.38 0.18 0.13 2.38 0.63 2.04 0.71 1.94 2.01 -1.99 -1.28
Convo Science 0.84 1.96 1.95 0.61 0.23 0.12 0.06 2.38 0.79 1.82 0.73 1.79 1.94 -2.23 -1.59
News+ Convo 0.82 1.34 1.53 0.40 0.19 0.11 -0.01 1.76 0.67 1.36 0.72 1.26 1.52 -5.29 -3.17
News+ Science 0.82 2.03 1.96 0.65 0.23 0.12 0.06 2.40 0.82 1.92 0.74 1.83 2.04 -2.40 -1.61
Science Convo 0.80 1.43 1.49 0.41 0.19 0.11 -0.02 1.77 0.64 1.34 0.67 1.28 1.58 -4.46 -2.61
Science News+ 0.78 1.96 2.07 0.44 0.38 0.18 0.13 2.33 0.62 1.95 0.68 1.94 1.99 -2.42 -1.51
SummN Convo News+ 0.90 2.71 2.64 0.64 0.23 0.14 0.01 2.75 0.89 2.92 0.82 2.74 2.18 -5.47 -8.28
Convo Science 0.77 2.51 2.36 0.72 0.21 0.13 0.01 2.48 0.92 2.61 0.88 2.27 1.99 -7.72 -11.28
News+ Convo 0.67 1.44 1.58 0.71 0.22 0.13 0.00 1.95 0.70 1.74 0.73 1.58 1.67 -3.45 -2.16
News+ Science 0.64 1.57 1.54 0.67 0.21 0.12 0.03 2.15 0.77 1.63 0.68 1.68 1.83 -1.39 -1.26
Science Convo 0.39 1.13 1.21 0.33 0.21 0.14 -0.04 1.26 0.63 1.37 0.33 1.25 1.26 -6.52 -4.14
Science News+ 0.65 1.95 2.04 0.33 0.31 0.17 0.04 2.36 0.69 2.05 0.35 2.07 1.90 -3.14 -3.58
Llama 3.1 Convo News+ 0.87 3.24 3.56 0.92 0.40 0.21 0.16 3.64 0.94 4.11 0.95 3.98 3.02 -2.31 -1.79
Convo Science 0.88 3.32 3.47 0.91 0.24 0.15 0.08 3.58 0.94 4.06 0.97 4.00 3.23 -1.39 -1.29
News+ Convo 0.82 3.01 3.23 0.94 0.22 0.15 0.05 3.43 0.94 3.99 0.96 3.92 2.96 -3.88 -1.36
News+ Science 0.88 3.38 3.45 0.92 0.24 0.15 0.09 3.61 0.93 4.03 0.98 4.00 3.19 -1.30 -1.26
Science Convo 0.82 3.17 3.21 0.94 0.22 0.15 0.06 3.47 0.93 3.94 0.96 3.92 2.91 -4.00 -1.46
Science News+ 0.87 3.25 3.54 0.94 0.40 0.22 0.16 3.64 0.95 4.07 0.95 3.98 2.93 -2.15 -1.69
Llama 3.1 (0-shot) zero-shot Convo 0.83 3.12 3.31 0.94 0.22 0.15 0.05 3.44 0.93 4.01 0.96 3.87 2.79 -3.74 -1.40
zero-shot News+ 0.88 3.26 3.41 0.93 0.41 0.22 0.16 3.57 0.94 4.04 0.94 4.00 3.02 -2.07 -1.64
zero-shot Science 0.87 3.42 3.52 0.93 0.24 0.16 0.08 3.47 0.93 4.06 0.98 3.96 3.29 -1.38 -1.30
@@
Table 4: Multi-document summarization (MDS) model performance when trained and tested on the different domains. Results are on the test splits of each dataset. For all metrics, a greater score is better.
Prometheus Evaluator Prompt
System You are a fair judge assistant tasked with providing clear, objective feedback based on specific criteria, ensuring each assessment reflects the absolute standards set for performance.
Assistant ###Task Description: An instruction (might include an Input inside it), a response to evaluate, a reference answer that gets a score of 5, and a score rubric representing a evaluation criteria are given. 1. Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general. 2. After writing a feedback, write a score that is an integer between 1 and 5. You should refer to the score rubric. 3. The output format should look as follows: "(write a feedback for criteria) [RESULT] (an integer number between 1 and 5)" 4. Please do not generate any other opening, closing, and explanations. ###The instruction to evaluate: You are a multi-document summarizer. When given a list of DOCUMENTS, you provide a SUMMARY that incorporates all of the documents. Write a summary that incorporates all of the following documents: ### DOCUMENTS {all source documents} ###Response to evaluate: {model-generated summary} ###Reference Answer (Score 5): {ground-truth summary} ###Score Rubrics: {rubric} ###Feedback:
Figure 9: The formatting is from Kim et al. (2024), and the instructions and rubric are from Figure˜10.
Evaluation Criteria for Prometheus
Factuality Is the model’s summary factually correct and well-supported by the DOCUMENTS?
Relevance How well does the summary capture the key points of the documents? Consider whether all and only the important aspects are contained in the summary. Only evaluate the summary for relevance. Do not penalize the summary for issues that are outside the scope of relevance.
Consistency How consistent are the facts in the summary with the facts in the original documents? Consider whether the summary does reproduce all facts accurately and does not make up untrue information. Do not use your own knowledge, and consider the documents as truth. Only evaluate the summary for consistency. Do not penalize the summary for issues that are outside the scope of consistency.
Fluency What is the quality of individual sentences, are they well-written and grammatically correct? Consider the quality of individual sentences.Only evaluate the summary for fluency. Do not penalize the summary for issues that are outside the scope of fluency.
Coherence What is the quality of all sentences collectively? How well do they fit together, and do they sound natural? Consider the quality of the summary as a whole. Only evaluate the summary for coherence. Do not penalize the summary for issues that are outside the scope of coherence.
Coverage Does the model’s summary incorporate salient information from all the documents? Do not penalize the summary for issues that are outside the scope of coverage.
Figure 10: Evaluation dimensions provided to Prometheus (Kim et al., 2024). The criteria for Relevance, Consistency, Fluency, and Coherence are drawn from SummEval (Fabbri et al., 2021b). We added the italicized text to the prompt to improve Prometheus performance.
Evaluation Criteria for UniEval (Summarization)
Relevance Is this summary relevant to the reference?
Consistency Is this claim consistent with the document?
Fluency Is this a fluent paragraph?
Coherence Is this a coherent summary to the document?
Figure 11: Evaluation dimensions for UniEval (Zhong et al., 2022). Only Relevance requires access to a reference summary.
Llama 3.1 Instruct Prompt Templates